ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 00, 01 Brecht, Gerhard, Fachhochschule Lippe, Lemgo, Germany, Oberschelp, Arnold, Christian-Albrechts-Universit“at zu Kiel, Kiel, Germany Do Mathematical Standards concern Mathematics? Mathematicians usually do not consider mathematical standards as necessary. Standards, while important for technology and trade, are considered to be superfluous in mathematics. It is often believed that mathematical standards might even disturb the creative unfolding of mathematical ideas. Of course, thinking about mathematical subjects is more important than thinking about mathematical terminology. But, good terminology will not in all cases evolve automatically. And users of mathematics (e. g. physicists and engineers) needing a unique and coherent mathematical terminology produce mathematical standards of their own (e. g. ISO 31-11). In the poster arguments are given why the question of mathematical standards should not be neglected by the mathematical community.